# Depth-Resolved Characterization of Centrifugal Disk Finishing of Additively Manufactured Inconel 718

### v1

@@ Abstract
Surface characteristics are a major contributor to the in-service performance, particularly fatigue life, of additively manufactured (AM) components. Centrifugal disk finishing (CDF) is one of many rigid media, abrasive machining processes employed to smooth the surfaces and edges of AM components. Within the general family of abrasive machining processes currently applied to AM, CDF is moderate in terms of material removal rate and the inertial forces exerted. How CDF alters the underlying microstructure of the processed surface is currently unknown. Here we employ profilometry and high-energy X-ray diffraction to characterize surface finish, crystallographic texture, and anisotropic distributions of residual microscale strain as a function of depth in CDF-finished Inconel 718 manufactured with laser powder bed fusion. Surfaces are finished using both unimodal and bimodal finishing media size distributions. We find that CDF will remove surface textures from AM components, but generally not alter the bulk texture. CDF is also found to impart significant amounts of residual microscale strain into the first 100 $\mu$m from the sample surface.
@@ 1 Introduction
Surface integrity is a controlling feature for the mechanical performance of metallic components, particularly the fatigue life. As-built surfaces of additively manufactured (AM) components often have poor surface integrity in the as-built state which is naturally detrimental to the mechanical response [1]. Primary machining processes are employed to subtract AM material to create functional surfaces with specific geometry and fine surface texture. For high performance components, secondary machining processes are additionally employed to subtract AM material from non-functional surfaces to smooth them. A common means to perform this secondary machining is to employ a family of processes referred to here as rigid media, abrasive machining processes. These processes are favored because they are capable of batch processing and are relatively inexpensive.
@@
Centrifugal disk finishing (CDF) is a mid-level, kinetic energy process that is commonly used for both semi-roughing and finishing of AM parts, with each process using a different media [2, 3]. The process employs a bowl with a rotating disk at the bottom. Media and parts are placed unrestrained in the bowl. The rotating disk energizes the media and parts, causing both to move rapidly through the bowl. While this occurs, an emulsion flows through the mix to lubricate the media and parts and to carry away fines. Media differ in shape, size, and composition to suit the geometry, starting surface texture, and hardness of the part.
@@
Although the effectiveness of rigid media, abrasive machining as a surface improvement approach is well established [2, 4] even in commercial settings, fundamental details of how the process alters the material microstructure below the surface is not well explored. For instance, Nalli et al. show that barrel finishing the surface of titanium alloy Ti6Al4V uniaxial tensile test specimens improves their yield strength, ultimate tensile strength and ductility [5], and Xiuhua et al. report increased surface hardness and compressive residual stresses in 18CrNiMo7-6 carburized rollers that underwent barrel finishing [6]. In fact, an increase in residual stress was also seen in Nickel super alloy Inconel 718 (IN718) that underwent barrel finishing [2]. However, these studies did not explore the microstructure-driven mechanisms that produced these effects. Understanding these effects can lead to optimization of process conditions beyond final surface condition and lead to superior mechanical performance.
@@
Optimization of abrasive treatments are complicated by process mechanics which often involves severe (sub-surface) plastic strains generated by stochastic interactions of the surface with the rigid abrasive media [7, 8]. While the surface finish may be improved [9], the microstructure and local residual strain and stress states below the surface are also altered which, in-turn, will alter the local mechanical response. Analysis of these sub-surface changes have traditionally been hampered by characterization limitations and as such are generally not optimized. Destructive serial-sectioning combined with microscopy will alter residual stress state, while low-energy, laboratory X-ray diffraction cannot penetrate metallic alloys with sufficient depth. Changes in microstructure can be inferred from surface hardness testing [10], but these measurements are not a direct measure of microstructural changes. However, synchrotron high-energy X-ray diffraction provides an ideal tool for non-destructively characterizing the effects of abrasive surface treatments on AM materials for process and microstructural optimization.
@@
Here we examine how variation of abrasive media and processing time alter surface state, along with depth-resolved subsurface microstructural and microstrain state. Characterization is performed using both optical profilometry and high-energy X-ray diffraction. We find clear variation in microscale stress state with varying processing route and alteration of surface preferred crystallographic orientation due to surface material removal. We begin by introducing the material and surface processing methods employed, followed by the characterization techniques. Results describe the final surface finish and sub-surface microstructures and microstrains. Implications for the application of centrifugal disk finishing to AM processes are then discussed.
@@ 2.1 Material and CDF Finishing
After removal, blocks were placed into a centrifugal finisher with triangular prism shaped ceramic (Al2O3) media obtained from Walther Trowal. Two media sizes were used with approximately 6 mm (WXC 6$\times$6 straight cut triangle, SCT) and 15 mm (WXC 15$\times$15 angle cut triangle, ACT) particle sizes. Four specimens were separately placed into a Walther Trowal TT45 centrifugal disc finisher under different surface processing conditions. Two specimens were processed in only the 6 mm particles (unimodal mixture) for 75 minutes and 135 minutes respectively, while another two specimens were processed in a 50:50 mixture of 6 mm and 15 mm particles (bimodal mixture) also for 75 minutes and 135 minutes respectively. One specimen was reserved with no surface finish (NF). After processing 0.9 mm thick specimens were extracted from the larger blocks. In total, 5 specimens were generated and the processing parameters are summarized in Table 1.
@@ 2.2 Surface Roughness Characterization
where $z_{i}$ is the height at each pixel in the surface roughness map, $\bar{z}$ is the mean value of $z_{i}$, and $n$ is the number of surface height measurements. Three positions were randomly chosen on each specimen, and the average of the surface roughness measures obtained from these spots are reported. The orientations these surface normals were always perpendicular to the build direction.
2.3 X-ray Characterization
@@
The diffraction images from the 5 specimens (3 volumes each) were processed using routines that take advantage of the HEXRD software package. Each diffraction image was broken into 36 10 azimuthal bins and then each bin was azimuthally (around the ring) integrated to produce one-dimensional intensity line profiles. For each line profile, Pseudo-Voigt peaks [12] were fit to the 111, 200, and 222 diffraction peaks (sets of lattice planes). From the fits, integrated intensity and full-width-at-half-maximum (FWHM) were extracted. The intensity and FWHM from the peaks were then mapped to 111, 200, and 222 pole figures in the sample coordinate system. From the rotation step size and azimuthal bin size, each pole figure has 648 measurement points. The intensity pole figure data were then transferred to the MTEX software [13] package for calculation of orientation distribution functions (ODFs). The ODF calculated $f$ follows the standard definition [14]
@@
where $\bm{r}$ is an orientation transformation from the local crystal to sample frame, $dV(\bm{r})$ is the volume fraction of crystal with orientation $\bm{r}$, and $V$ is the diffraction volume. Scalar texture metrics are also calculated using MTEX including the texture strength $T$ to quantify the disorder of the texture defined as
@@
and the entropy $S$ to quantify defined as
@@
Lastly, the fit ODFs were re-projected to pole figures for texture data presentation.
@@ 3.1 Surface Measurements
The surface in the NF state exhibited a rough texture that featured roughness measures of $R_{a}\sim 4.9\ \mu$m, $R_{q}\sim 6.5\ \mu$m, and $R_{z}\sim 52.7\ \mu$m. An example of the texture resulting from AM is shown in Fig. 2a. As expected, this NF texture comprised waviness that often results from the process of additive (layer-by-layer) fabrication. In addition, the figure also shows randomly positioned peaks each spanning a much smaller length scale of approximately 20 to 80 $\mu$m diameter. These peaks originated from partially melted powder particles that adhere to the surface during AM.
Figure 2: Surface topographies of representative specimens in the following stages of CDF: a) NF (not finished), b) U1 (6$\times$ 6 SCT media, 75 mins), c) U2 (6$\times$ 6 SCT media, 135 mins), d) B1 (bimodal media, 75 mins), e) (bimodal media, 135 mins).
@@ 3.2 X-ray Measurements
As described, the textures of the samples were measured at three depths away from the surface. Pole figures from the 111, 200, and 220 lattice planes in the NF specimen are projected from the ODFs calculated and presented in Fig. 3 as a function of depth: 20 $\mu$m (Fig. 3a), 70 $\mu$m (Fig. 3b), and 120 $\mu$m (Fig. 3c). The orientation pole figure data is presented in multiples of uniform density (MUD). In Fig. 3, we see that the NF specimen has a relatively strong fiber texture at the sample surface (20 $\mu$m depth) with the {111} lattice plane normals nominally aligned with the build direction and the sample surface. The grains in this diffraction volume are randomly oriented about the [111] direction as evidenced by the rings of intensity observed in the 200 and 220 pole figures. With increasing depth, the texture transitions from the 111 fiber texture to a weak cube texture which we can see from the larger MUD values on the 200 pole figures along the $\bm{y^{S}}$ (build) direction and the 4-fold symmetry of intensity about $\bm{y^{S}}$ on all three pole figures.
@@
Figure 3: Comparisons of the measured textures in the NF (no finish) specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Similar to the NF sample, texture data for the U1 and B1 specimens (see Table 1) are shown in 4 and Fig. 5 respectively. The textures for the U2 and B2 specimens are not presented as they are qualitatively similar to U1 and B1. All measured textures with depth in both U1 and B1 specimens (Fig. 4 and Fig. 5 respectively) are relatively similar. Each exhibit a cube texture similar to the lowest depth (120 $\mu$m) in Fig. 3c, but with higher peak values. In both samples, the peak intensity values in the pole figures appear to decrease with decreasing depth, implying the surface treatments reinforce the existing cube texture from the AM build. The lack of the 111 fiber texture in the CDF-finished samples indicates that approximately 100 $\mu$m of material was removed from the samples and along with it, the remnant surface texture from the AM build process.
@@
Figure 4: Comparisons of the measured textures in the U1 (6$\times$ 6 SCT media, 75 mins) surface-finished specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Figure 5: Comparisons of the measured textures in the B1 (bimodal media, 75 mins) surface-finished specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Figure 6a shows the variation of the ODF strength $T$ with depth. We can see that in the NF specimen, the strongest texture is at the sample surface which decreases with depth, eventually reaching texture strengths comparable to the surface-finished specimens. This is consistent with the abrasive media removing the top surface layer of the AM build. In the surface finished specimens, the texture strengths are comparable, with the B1 texture being the strongest. However, the relatively constant texture strength with depth implies that the increased texture strength B1 is most likely a remnant of the initial texture of the AM build. The texture strengths appear to peak at the 70 $\mu$m depth, in all specimens but this increase is small and within measurement uncertainty. Similar findings, but reversed, can be seen in Fig. 6b showing the variation ODF entropy $S$ with depth.
@@
Figure 6: a) Variation of the ODF strength $T$ in the five specimens as a function of depth. b) Variation of the ODF strength $S$ as a function of depth.
@@
To examine the microscale residual strains present in the material, the full-width-at-half-maximum (FWHM) of the various sets of lattice planes with respect to sample direction were also extracted from the diffraction data. The FWHM of a diffraction peak captures the spread of lattice plane spacing across a diffraction volume which will generally have contributions from stress features across length scales. In this case, the peak FWHM primarily has contributions from residual stresses generated during grain-scale plastic deformation and local defect (dislocation) content. We note that as the samples were textured, FWHM values from diffraction peaks with negligible intensity are omitted.

### v2

@@ Abstract
Surface characteristics are a major contributor to the in-service performance, particularly fatigue life, of additively manufactured (AM) components. Centrifugal disk finishing (CDF) is one of many rigid media, abrasive machining processes employed to smooth the surfaces and edges of AM components. Within the general family of abrasive machining processes currently applied to AM, CDF is moderate in terms of material removal rate and the inertial forces exerted. How CDF alters the underlying microstructure of the processed surface is currently unknown. Here we employ white light profilometry and high-energy X-ray diffraction to characterize surface finish, crystallographic texture, and anisotropic distributions of residual microscale strain as a function of depth in CDF-finished Inconel 718 manufactured with laser powder bed fusion. Surfaces are finished using both unimodal and bimodal finishing media size distributions. We find that CDF will remove surface crystallographic textures (here a {111} fiber texture) from AM components, but generally not alter the bulk texture (here a cube texture). CDF is also found to impart significant amounts of residual microscale strain into the first 100 $\mu$m from the sample surface as evidenced by an approximately 50% increase in diffraction peak widths at 20 $\mu$m from the surface in comparison to 120 $\mu$m.
@@ 1 Introduction
Surface integrity is a controlling feature for the mechanical performance of metallic components, particularly the fatigue life. Surfaces of additively manufactured (AM) components often have poor surface integrity in the as-built state which is naturally detrimental to the mechanical response [1, 2]. Primary machining processes are employed to subtract AM material to create functional surfaces with specific geometry and fine surface texture. For high performance components, secondary machining processes are additionally employed to subtract AM material from non-functional surfaces to smooth them. A common means to perform this secondary machining is to employ a family of processes referred to here as rigid media, abrasive machining processes. These processes are favored because they are capable of batch processing and are relatively inexpensive.
@@
Centrifugal disk finishing (CDF) is a mid-level, kinetic energy process that is commonly used for both semi-roughing and finishing of AM parts, with each process using a different media [3, 4]. The process employs a bowl with a rotating disk at the bottom which is illustrated schematically in Fig. 1. Media and parts are placed unrestrained in the bowl. The rotating disk energizes the media and parts, causing both to move rapidly through the bowl. While this occurs, an emulsion flows through the mix to lubricate the media and parts and to carry away fines. Media differ in shape, size, and composition to suit the geometry, starting surface texture, and hardness of the part. The smoothing effect in CDF is shear-based and originates from mechanical interaction between the media used and the surface. This mechanical interaction can also alter the microstructure and the stress state of the material in the vicinity of the surface of the part. In this regard, the evolution of surface integrity of a part during CDF, characterized by its surface roughness and crystallographic texture is coupled.
Figure 1: A schematic of the centrifugal disk finishing process in which abrasive media and components are placed into a cylindrical bowl. A rotating disc at the bottom of the assembly excites media and component motion to induce finishing.
@@
Although the effectiveness of rigid media, abrasive machining as a surface improvement approach is well established [5, 3, 6] even in commercial settings, fundamental details of how the process alters the material microstructure below the surface is not well explored. For instance, Nalli et al. show that barrel finishing the surface of titanium alloy Ti6Al4V uniaxial tensile test specimens improves their yield strength, ultimate tensile strength and ductility [7], but Denti et al. found that tumbling alone was not sufficient to improve fatigue life of laser powder bed fusion Ti6Al4V and subsequent shot-peening was required. [8]. Xiuhua et al. report increased surface hardness and compressive residual stresses in 18CrNiMo7-6 carburized rollers that underwent barrel finishing [9]. In fact, an increase in residual stress was also seen in nickel super alloy Inconel 718 that underwent barrel finishing [3]. Together these results imply that improving surface finish alone is not sufficient to control final mechanical performance of AM parts, and the microstructure and residual stress modification also play a role. Understanding these effects can lead to optimization of process conditions beyond final surface condition and lead to superior mechanical performance.
@@
Optimization of abrasive treatments are complicated by process mechanics which often involve severe (subsurface) plastic strains generated by stochastic interactions of the surface with the rigid abrasive media [10, 11]. While the surface finish may be improved [12], the microstructure and local residual strain and stress states below the surface are also altered which, in-turn, will alter the local mechanical response. Analysis of these subsurface changes have traditionally been hampered by characterization limitations and as such are generally not optimized. Destructive serial-sectioning combined with microscopy will alter residual stress state, while low-energy, laboratory X-ray diffraction cannot penetrate metallic alloys with sufficient depth. Changes in microstructure can be inferred from surface hardness testing [13], but these measurements are not a direct measure of microstructural changes. However, synchrotron high-energy X-ray diffraction provides an ideal tool for non-destructively characterizing the effects of abrasive surface treatments on AM materials for process and microstructural optimization.
Here we examine how variation of abrasive media and processing time alter surface state, along with depth-resolved subsurface microstructural and microstrain state in AM Inconel 718 (IN718) built using laser powder bed fusion. IN718 nickel superalloy is generally employed for its high-temperature strength and corrosion resistance [14, 15] and its machining [16, 17, 18, 19, 20] and finishing [21, 22, 23] response has been extensively studied. This material, along with Inconel 625, has also been a focus of intense study by the additive manufacturing community [24, 25, 26]. Characterization is performed using both white light profilometry and high-energy X-ray diffraction. We find clear variation in microscale stress state with varying processing route and alteration of surface preferred crystallographic orientation due to surface material removal. In addition, we show that as surface roughness decreases, the magnitude of residual stresses increases during CDF. However, the mechanical interactions between the media and the surface that produced this effect were not strong enough to cause evolution of crystallographic textures in the material close to the surface. This implies a crystallographic texture-preservation effect, which is unexpected in shear-based finishing processes [11, 27]. This is an important result as the surface integrity of a part, which is characterized by its roughness, residual stress, and crystallographic texture, significantly influences its ability to resist damage during fatigue. Herein, the ability to preserve beneficial crystallographic textures near the surface can provide novel avenues for enhancing a part’s life, particularly during cyclic loading conditions.
@@
This work is structured as follows. We begin by introducing the material and surface processing methods employed, followed by the characterization techniques. Results describe the final surface finish and subsurface microstructures and microstrains. Implications for the application of centrifugal disk finishing to AM processes are then discussed.
@@ 2.1 Material and CDF Finishing
After removal, blocks were placed into a centrifugal finisher with triangular prism shaped ceramic (Al2O3) media obtained from Walther Trowal. Two media sizes were used with approximately 6 mm (WXC 6$\times$6 straight cut triangle, SCT) and 15 mm (WXC 15$\times$15 angle cut triangle, ACT) particle sizes. Four specimens were separately placed into a Walther Trowal TT45 centrifugal disc finisher under different surface processing conditions. Finishing was carried out in a lubricated condition. The KFL compound purchased from Walther Trowal was used as a lubricant and endowed to the media at the rate 300 mL/min. The speed of revolution for finishing was 250 rpm for all the specimens. A cumulative total amount of 25 kg of media was used for both, unimodal and bimodal conditions. Two specimens were processed in only the 6 mm particles (unimodal mixture) for 75 minutes and 135 minutes respectively (U1 and U2), while another two specimens were processed in a 50:50 mixture of 6 mm and 15 mm particles (bimodal mixture) also for 75 minutes and 135 minutes respectively (B1 and B2). One specimen was reserved with no surface finish (NF). After processing 0.9 mm thick specimens were extracted from the larger blocks. In total, 5 specimens were generated and the processing parameters are summarized in Table 1. A photo of the 5 specimens is provided in Fig. 2 where the finishing effect of the CDF process can be clearly visualized between the NF and finished specimens.
@@
Figure 2: A photo of the 5 surface-finish specimens characterized with high-energy X-ray diffraction. a) NF (not finished), b) U1 (6$\times$6 SCT media, 75 mins), c) U2 (6$\times$6 SCT media, 135 mins), d) B1 (bimodal media, 75 mins), e) B2 (bimodal media, 135 mins).
@@ 2.2 Surface Roughness Characterization
where $z_{i}$ is the height at each pixel in the surface roughness map, $\bar{z}$ is the mean value of $z_{i}$, and $n$ is the number of surface height measurements. Three positions were randomly chosen on each specimen, and the average of the surface roughness measures obtained from these spots are reported. The orientations of these surface normals were always perpendicular to the build direction.
The choice of these parameters for the current study was motivated by their widespread use in industry. Among these, benchmark values for the $R_{a}$ parameter for various applications such as sealing, and fatigue loading are well known. This makes it easier to qualify a part for a certain application, e.g., if its $R_{a}$ value is smaller than the pertinent benchmark. Nonetheless, these parameters provide an incomplete picture of the surface [13]. More sophisticated parameters that have a systematically calculable physical meaning, e.g., surface curvature, can be used to achieve a more comprehensive description of the condition of the surface [28]. However, these parameters are more challenging to calculate which limits their widespread adoption. In the discussion, a more complex measure of surface roughness, entropy, is used to understand the mechanics of surface evolution during CDF.
2.3 X-ray Characterization and Data Processing
@@
The diffraction images from the 5 specimens (3 volumes each) were processed using routines that take advantage of the HEXRD software package [30, 31]. Each diffraction image was broken into 36 10 azimuthal bins and then each bin was azimuthally (around the ring) integrated to produce one-dimensional intensity line profiles. The azimuthal integration process consists of remapping each diffraction image to a polar coordinate system described by radial and azimuthal directions centered around the intersection of the incoming beam with the detector using bi-linear interpolation. The polar-remapped diffracted intensity is then integrated by summing intensity along the azimuthal direction. For each line profile, Pseudo-Voigt peaks [32] (Gaussian and Lorentzian functions combined with a rule of mixtures) were fit to the 111, 200, and 222 diffraction peaks (sets of lattice planes). From the fits, integrated intensity and full-width-at-half-maximum (FWHM) were extracted. The intensity and FWHM from the diffraction peaks were then mapped to 111, 200, and 222 pole figures in the sample coordinate system. From the rotation step size and azimuthal bin size, each pole figure has 648 measurement points. The intensity pole figure data were then transferred to the MTEX software package [33] for calculation of orientation distribution functions (ODFs). The ODF calculated $f$ follows the standard definition [34]
@@
where $\bm{r}$ is an orientation transformation from the local crystal to sample frame, $dV(\bm{r})$ is the volume fraction of crystal with orientation $\bm{r}$, and $V$ is the diffraction volume. ODFs are determined by minimizing the difference (L2 norm) between projections of the fit ODF onto synthetic pole figures and the measured pole figure data. Scalar crystallographic texture metrics are also calculated using MTEX including the strength $T$ of the crystallographic texture defined as
@@
and the entropy $S$, to quantify the disorder, defined as
@@
Lastly, the fit ODFs were re-projected to pole figures for crystallographic texture data presentation.
@@ 3.1 Surface Measurements
The surface in the NF state exhibited a rough texture that featured roughness measures of $R_{a}\sim 4.9\ \mu$m, $R_{q}\sim 6.5\ \mu$m, and $R_{z}\sim 52.7\ \mu$m. An example of the texture resulting from AM is shown in Fig. 4a. As expected, this NF texture contains waviness that often results from the process of additive (layer-by-layer) fabrication. In addition, the figure also shows randomly positioned peaks each spanning a much smaller length scale of approximately 20 to 80 $\mu$m diameter. These peaks originated from partially melted powder particles that adhere to the surface during AM.
Figure 4: Surface topographies of representative specimens in the following stages of CDF: a) NF (not finished), b) U1 (6$\times$6 SCT media, 75 mins), c) U2 (6$\times$6 SCT media, 135 mins), d) B1 (bimodal media, 75 mins), e) B2 (bimodal media, 135 mins).
@@
The repeatability of the measurements obtained using white light profilometry is provided in the error bars (2 $\times$ the error $e$) associated to the quantified surface roughness metrics ($R+e$) which are reported in Table 2. These error bars were characterized from 3 different measurements that were made on 3 different randomly chosen points on every specimen under equivalent conditions. A maximum error of $1.302e=$ 1.30 $\mu$m was determined for surface roughness metric $R_{a}$ that was obtained from the U2 specimen prior to finishing. We note that this value is significantly larger than the resolution of the optical profilometer (reported as several angstroms) and is likely a consequence of how rough the surfaces themselves were in this state.
We anticipate that the error associated with surface roughness metrics would inflate if sufficiently large areas were not measured during white light profilometry. Nonetheless, increasing the dimensions of the measured regions also comes at the cost of decreasing spatial resolution of the measured surface topographies. To this end, we note that the dimensions of our measured regions ($\sim$ 0.85 mm $\times$ 0.85 mm) were similar to those in other well-established works within this focus area ($\sim$ 1.5 mm $\times$ 1 mm) [35].
@@ 3.2 X-ray Measurements
As described, the crystallographic textures of the samples were measured at three depths away from the surface. Pole figures from the 111, 200, and 220 lattice planes in the NF specimen are projected from the ODFs calculated and presented in Fig. 5 as a function of depth: 20 $\mu$m (Fig. 5a), 70 $\mu$m (Fig. 5b), and 120 $\mu$m (Fig. 5c). The orientation pole figure data is presented in multiples of uniform density (MUD). In Fig. 5, we see that the NF specimen has a relatively strong fiber texture at the sample surface (20 $\mu$m depth) with the {111} lattice plane normals nominally aligned with the build direction and the sample surface. The grains in this diffraction volume are randomly oriented about the [111] direction as evidenced by the rings of intensity observed in the 200 and 220 pole figures. With increasing depth, the texture transitions from the 111 fiber texture to a weak cube texture which we can see from the larger MUD values on the 200 pole figures along the $\bm{y^{S}}$ (build) direction and the 4-fold symmetry of intensity about $\bm{y^{S}}$ on all three pole figures.
@@
Figure 5: Comparisons of the measured crystallographic textures in the NF (no finish) specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Similar to the NF sample, texture data for the U1 and B1 specimens (see Table 1) are shown in 6 and Fig. 7 respectively. The textures for the U2 and B2 specimens are not presented as they are qualitatively similar to U1 and B1. All measured textures with depth in both U1 and B1 specimens (Fig. 6 and Fig. 7 respectively) are relatively similar. Each exhibit a cube texture similar to the lowest depth (120 $\mu$m) in Fig. 5c, but with higher peak values. In both samples, the peak intensity values in the pole figures appear to decrease with decreasing depth, implying the surface treatments reinforce the existing cube texture from the AM build. The lack of the 111 fiber texture in the CDF-finished samples indicates that approximately 100 $\mu$m of material was removed from the samples and along with it, the remnant surface texture from the AM build process.
@@
Figure 6: Comparisons of the measured crystallographic textures in the U1 (6$\times$6 SCT media, 75 mins) surface-finished specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Figure 7: Comparisons of the measured crystallographic textures in the B1 (bimodal media, 75 mins) surface-finished specimen at various depths: a) 20 $\mu$m, b) 70 $\mu$m, and c) 120 $\mu$m.
@@
Figure 8a shows the variation of the ODF strength $T$ with depth. We can see that in the NF specimen, the strongest crystallographic texture is at the sample surface which decreases with depth, eventually reaching crystallographic texture strengths comparable to the surface-finished specimens. This is consistent with the abrasive media removing the top surface layer of the AM build. In the surface finished specimens, the crystallographic texture strengths are comparable, with the B1 texture being the strongest. However, the relatively constant crystallographic texture strength with depth implies that the increased strength B1 is most likely a remnant of the initial crystallographic texture of the AM build. The crystallographic texture strengths appear to peak at the 70 $\mu$m depth, in all specimens but this increase is small and within measurement uncertainty. Similar findings, but reversed, can be seen in Fig. 8b showing the variation ODF entropy $S$ with depth.
@@
Figure 8: a) Variation of the ODF strength $T$ in the five specimens as a function of depth. b) Variation of the ODF entropy $S$ as a function of depth.
@@
To examine the microscale residual strains present in the material, the full-width-at-half-maximum (FWHM) of the various sets of lattice planes with respect to sample direction were also extracted from the diffraction data. The FWHM of a diffraction peak captures the spread of lattice plane spacing across a diffraction volume which will generally have contributions from stress features across length scales. In this case, the peak FWHM primarily has contributions from residual stresses generated during grain-scale plastic deformation and local defect (dislocation) content. We note that as the samples were crystallographic textured, FWHM values from diffraction peaks with negligible intensity are omitted.